OpenStudy (anonymous):

what is the exact value of csc 135

OpenStudy (jdoe0001):

check your Unit Circle, recall that \(\bf csc(\theta)=\cfrac{1}{sin(\theta)}\)

OpenStudy (sweetburger):

are you looking for like a pi/radian value?

OpenStudy (solomonzelman):

Yes, or use sin(a+b)=sin(a)cos(b)+sin(b)cos(a) \(\large\color{blue}{ \rm sin(135)=sin(90+45)=....}\) and then set the result over 1.

OpenStudy (anonymous):

guys i am still confused..

OpenStudy (anonymous):

nvm, got it